# Shared Control with Black Box Agents using Oracle Queries

### v1

@@ Abstract
Shared control problems involve a robot learning to collaborate with a human. When learning a shared control policy, short communication between the agents can often significantly reduce running times and improve the system’s accuracy. We extend the shared control problem to include the ability to directly query a cooperating agent. We consider two types of potential responses to a query, namely oracles: one that can provide the learner with the best action they should take, even when that action might be myopically wrong, and one with a bounded knowledge limited to its part of the system. Given this additional information channel, this work further presents three heuristics for choosing when to query: reinforcement learning-based, utility-based, and entropy-based. These heuristics aim to reduce a system’s overall learning cost. Empirical results on two environments show the benefits of querying to learn a better control policy and the tradeoffs between the proposed heuristics.
1 Introduction
Many modern systems require a human and a robotic system to work jointly in a coordinated manner. Surgery and service robots [1, 2], semi-autonomous vehicles [3, 4], and brain-computer interfaces [5] all involve combining the actions instructed by a human and a robot to form a shared control system. A shared control, according to Abbink et al. [6], is a system where “Human(s) and robot(s) are interacting congruently in a perception-action cycle to perform a dynamic task that either of them could execute individually under ideal circumstances.” A key challenge in designing such a system is that the policy of the human agent is hidden from the artificial agent. The human’s hidden policy can be viewed as a black box. However, existing solutions often assume learners can only interact with the black box via execution [7, 8]. In many of the mentioned applications, the black box is a cooperative agent, so a simple question can significantly improve the learning speed and accuracy of the shared system. This work aims to study the use of queries to improve learning in a shared control system. This goal is achieved by proposing a new framework for shared control as visualized in Figure 1. The framework comprises two systems: a black box (representing the human operator) and a control system (representing the robot).
In the proposed framework, the control can use the assistance of an oracle, which will restrict its potential action space. With the help of such an oracle, learning an efficient policy is expected to be easier, which in turn will minimize the number of failed operations or interactions in the shared system. In this work, we use two oracle types to demonstrate that this expectation is met when the oracle has full knowledge of the shared system, but in other cases, its advice can actually hinder the system’s performance. Once queries become part of the robot’s optional learning process, a key challenge becomes reasoning about when to query the oracle. This paper presents three heuristics that are designed to increase the accuracy of the learned control policy and to minimize the number of queries used during training and execution.
The contributions of this work are thus threefold: (1) A formal definition of the shared control problem with queries; (2) An introduction of two oracle types; and (3) Three heuristics for querying. These contributions are evaluated on two domains: the first is a set of carefully curated shared automata that magnify the unique challenges of learning a shared control policy, and the second is a commonly used simulation of a lunar lander operation [9] to show the heuristics’ performance in a more complex environment. The results show that using queries in those domains can significantly improve learning (even when the oracle’s suggestion might be misleading).
2 Related Work
@@
Given a shared control system, a key challenge is to design or learn a control that can work with the black box agent to maximize the return of the shared execution. Dragan and Srinivasa [7] proposed to view the shared control problem as an inverse reinforcement learning (IRL) problem, in which the task is to elicit the perceived reward of humans from their joint interactions with the control. This model was recently extended to Bayesian IRL [10]. Alternatively, Flad et al. [3] presented a system that models a human driver’s steering motion as a sequence of motion primitives and then uses these primitives to predict the resulting torque. Jiang et al. [11] model a higher-level layer of the intentions of the human rather than their motor primitives. Dann et al. [12] predict other agents’ intentions, assuming all agents have a similar underlying model.Iosti et al. [8] learns a shared control for black box environments using a Recurrent Neural Network (RNN). We use this latter solution as the underlying learner for the control agent, as it is the most unstructured solution, not modeling reward or goal estimations but only the state and actions.
@@
As often the black box agent in the shared control system is actually a human or a cooperating agent [13, 14, 15, 16, 17], we wish to leverage an additional communication channel, where the black box assumption is relaxed to allow the control to query the black box about the next best action it should take. Querying in such a manner to improve learning has been used in a variety of learning and inference problems, including active machine learning [18, 19, 20], diagnosis [21, 22], robotics [23, 24, 25] and planning [26, 27, 28]. However, to the best of our knowledge, this work is the first to leverage active learning in shared control systems. We formalize this new communication channel by incorporating an oracle into the shared control system.
3 The Query-based Shared Control Problem
@@
We start by formalizing the query-based shared control problem using a Multi-Agent Markov Decision Process (MA-MDP) that combines the state and action spaces of the black box agent and the control [29]. The key uniqueness of this model is in splitting the state space into two, according to the visibility of the agents’ state, from the learning agent’s perspective. This separation of the state space can also be viewed as an extension of Hidden-mode MDP into multiagent settings [30], and it differentiates from a decentralized POMDP formulation (DEC-POMDP) as the model and distribution over actions of other agents is unknown [31]. Each of the two spaces can capture an aggregated state of more than one agent in the environment and thus model systems with more than two agents. However, for brevity, in this paper we focus on the simplest case where the hidden part of the state space represents the state of a single black-box agent ($\mathbb{E}$), and the visible part represents the state space of the control ($\mathbb{C}$).
@@
The objective of the shared control can be defined using $R$ to maximize different objectives, such as reaching some goal (as in shared driving, like Example 1), maximizing the return (as in the Lunar Lander domain) or synchronizing the actions of the components (as in the Automata domain). In this paper, we keep the transition function deterministic, but it can straightforwardly be extended to stochastic transitions. The transition will be determined according to an Operation Protocol (OP), defining the way that the actions of the agents are combined.
@@
The proposed definition of a shared control system is very broad and can encompass a variety of shared control protocols. Different protocols can let one agent or both manipulate the joint operation or override an action. Alternatively, if the state and action spaces of the control and the black box are the same, a shared control system can blend the chosen actions into some weighted sum.
To enable communication between a human and a robot, we define an oracle to be a function that gets a joint state and returns a set of actions $O:S^{\mathbb{C}}\times S^{\mathbb{E}}\rightarrow\mathcal{P}(A^{\mathbb{C}}\cup A^{\mathbb{E}})$. We do not restrict the answer of the oracle to be an action that is executable by the control or the black box — the answer depends on the type of oracle used. E.g. a teacher oracle is familiar with the system as a whole and can provide the best action(s) the control should take to accomplish the shared objective over the course of an entire execution: $O_{Teacher}(s^{\mathbb{C}},s^{\mathbb{E}})=\{\operatorname*{arg\,max}_{a\in A^{\mathbb{C}}}G(s^{\mathbb{C}},a)\}$, Where $G(s^{\mathbb{C}},a)$ can lead to the best potential return if the control takes action $a$ in $s^{\mathbb{C}}$. An expert oracle is one that is familiar with the black box’s behavior and can predict what is the next action that will maximize its reward: $O_{Expert}(s^{\mathbb{C}},s^{\mathbb{E}})=\{\operatorname*{arg\,max}_{a\in A^{\mathbb{E}}}R(s^{\mathbb{E}},a)\}$.
@@
4 Query Heuristics for a Shared Control
To answer the question of when to query the oracle, we propose three different heuristics, each inspired by different factors: The Entropy-based heuristic takes an information-theoretical perspective; A Utility-based heuristic is based on prior knowledge; and an RL-based heuristic.
Entropy Heuristic The gist of this heuristic is to reason about the confidence of the learner about the information that was collected so far (or whether there’s a need to use an oracle). Information Gain (IG) is a measurement that can provide an answer to this question. When considering the IG with respect to an oracle, $S$ is the state of the joint system and $q$ is the query. Then, $H(S)$ is the entropy when acting without querying, and similarly, $H(S\mid q)$ is the conditional entropy of $S$ given that we query the oracle. When the IG is bigger than a certain threshold, the system will use the aid of an oracle to choose the next step. A key part of using the entropy heuristic to decide when to query is computing the IG. Given a shared control system $M$, $a^{\mathbb{C}}\in A^{\mathbb{C}}$ the action that was chosen by the control, $a^{\mathbb{E}}\in A^{\mathbb{E}}$ the action that was chosen by the black box, and $a^{o}\in A^{\mathbb{C}}\cup A^{\mathbb{E}}$ the action that was chosen by the oracle. The IG of the oracle for a state $s\in S^{\mathbb{C}}\times S^{\mathbb{E}}$ is the difference between querying or not: $IG(s,querying)=H(s)-H(s\mid querying)$, where
@@ 4 Query Heuristics for a Shared Control
Where $H(s,a^{o})$ represents the entropy of the control’s choices when using the oracle, and receiving from it action $a^{o}$ in a state $s$. To conclude, given a state $s$ and a threshold $\beta_{ent}$, this heuristic chooses to query ($Oracle(s)$) or use its learned action ($Learned(s)$) according to the following condition:
@@
Utility Heuristic This heuristic adds a preliminary step before choosing whether to query or not. When in state $s$, it first checks what actions are available to the control in $s$, and uses the suggestion of the RNN if there is an action with a probability higher than some chosen threshold $\beta_{util}$.
@@
Reinforcement Learning Heuristic An RL problem is a tuple $\langle S,A,R\rangle$ where $S$ is a set of states, $A$ is a set of actions the agent can execute in the environment, and $R:S\times A\rightarrow\mathbb{R}$ is the reward for executing $a$ at a state $s$ (we do not model the transition function as it is unknown to the control). In the shared control learning context, the state space used in this heuristic is the states of the visible control $S^{\mathbb{C}}$, the action space consists of two actions: to query and not to query, and the reward is defined to some positive reward if after a choice whether to query or not to query. Notice that this MDP is different from the MA-MDP used to model the shared control system as a whole. The key difference is that here, the agent’s decision is only about querying. The RL heuristic is based on the Q-learning algorithm [32], where the agent keeps an estimation of the value of executing an action in a state using a value function. This function is called a Q-function, defined as $Q:S\rightarrow A$. The heuristic chooses to query if the value of querying is higher than not.
@@
5 Expeiments
The underlying learning agent used in this work is the one presented in Iosti et al. [8]. However, as mentioned earlier, it can be replaced by any other learning agent that can process the states and actions of the system and recommend the next action of the control. The evaluation in this paper consists of two domains: an automata-based system and an Atari-based lunar lander simulation [33].
Domain 1: Automata-based Shared Control We follow the definition of an automaton by Iosti et al. [8]. These automata are carefully chosen to challenge the learning of the control policy, due to sparse or delayed rewards. An automaton A is a tuple $\Omega=\langle S,\iota,A,\delta\rangle$, where $S$ is a finite set of states, $\iota\in S$ it the initial state of the automaton, $A$ is a finite set of actions, and $\delta:(S\times A)\rightarrow S\cup\{\perp\}$ is a partial transition function, where $\perp$ stands for an undefined transition. For brevity, we use $a(s)$ to define the transition $\delta(a,s)$. Denoting $en(s)=\{a\mid a\in A\wedge\delta(s,a)\neq\perp\}$, i.e., $en(s)$ is the set of actions enabled at the state $s$. We assume that for each $s\in S$, the set of possible actions is not empty, $en(s)\neq\emptyset$. A shared control system consists of two automata, a black-box ($\Omega^{\mathbb{E}}$) and the control ($\Omega^{\mathbb{C}}$). As this is a black-box environment, the control automaton is visible to the learner, but the automaton is not. The operation protocol for the two automata is an agreement operator, which is defined as:

### v2

@@ Abstract
Shared control problems deal with an agent learning to act in collaboration with other agents or systems. When learning a shared control policy, often a short communication between the agents can significantly reduce running times and improve the system’s accuracy. We extend the shared control problem to include the ability to directly query a cooperating agent. We consider two types of potential responses to a query, namely oracles: one that assumes that the oracle has a perfect knowledge of the shared system and can provide the learner with the best action it should take, even when that action might be myopically wrong (a teacher), and one with a bounded knowledge limited to its own part of the system (an expert). Given this additional information channel, this work presents several heuristics for choosing when to query: reinforcement learning-based, utility-based, and entropy-based. These heuristics are aimed at reducing the overall learning cost of a system. Empirical results on two domains show the benefits of querying to learn a better control policy and demonstrate the tradeoffs between the proposed heuristics.
I Introduction
Many modern systems require a human and an artificial agent to work jointly in a coordinated manner. Surgery and service robots [23, 29], autonomous vehicles [11, 17], and brain-computer interfaces [22] all involve combining the actions instructed by a human and an artificial agent to form a shared control system. According to Abbink et al. [1],
  • “In shared control, human(s) and robot(s) are interacting congruently in a perception-action cycle to perform a dynamic task, that either of them could execute individually under ideal circumstances.”
A key challenge in the design of such a system is that the policy of the human agent is hidden from the artificial agent. The human’s hidden policy can be viewed as a black box. However, existing solutions often assume that the learner can only interact with the black box via execution [9, 12]. In many of the mentioned applications, the black box is a cooperative agent, so a simple question can result in great improvements to the learning speed and accuracy of the shared system.
The goal of this work is to study the use of queries to improve learning in a shared control system. This goal is achieved by proposing a new framework for shared control as visualized in Figure 1. The framework comprises two systems: a black box (representing the human operator) and a control system (representing the autonomous agent).
@@ I Introduction
In the proposed framework, the control can use the assistance of an oracle, which will restrict its potential action space. The oracle may be a cooperating human or an external expert. With the help of such an oracle, learning an efficient policy is expected to be easier, which in turn will minimize the number of failed operations or interactions in the shared system. In this work, we use two oracles types to demonstrate that this expectation is met when the oracle has full knowledge of the shared system, but in other cases, its advice can actually hinder the system’s performance. Once queries become part of the autonomous agent’s optional learning process, a key challenge becomes reasoning about when to query the oracle. This paper presents three heuristics that are designed to increase the accuracy of the learned control policy and to minimize the number of queries used during training and execution.
The contributions of this work are thus threefold: (1) A formal definition of the shared control problem when augmented with queries; (2) An introduction of two oracle types for queries; and (3) Three heuristics for querying in shared control systems. These contributions are evaluated on two domains: the first is a set of carefully curated shared automata that magnify the unique challenges of learning a shared control policy, and the second is a commonly used shared control simulation of a lunar lander operation [6] to show the heuristics’ performance in a more natural environment. The results show that the use of queries in those domains can significantly improve learning in both domains (even when the oracle’s suggestion might be misleading) and that the different heuristics offer a trade-off between the number of queries used and the accuracy of the learned policy.
II Related Work
@@
Given a shared control system, a key challenge is to design or learn a control that can work with the black box agent to maximize the return of the shared execution. Dragan and Srinivasa [9] proposed to view the shared control problem as an inverse reinforcement learning problem, in which the task is to elicit the perceived reward of humans from their joint interactions with the control. Alternatively, Flad et al. [11] presented a system that models a human driver’s steering motion as a sequence of motion primitives and then uses these primitives to predict the resulting torque. Jiang et al. [15] model a higher-level layer of the intentions of the human rather than their motor primitives. Then, they predict how the shared system should act to accomplish those goals. Iosti et al. [12] learn a shared control for black box environments using a Recurrent Neural Network (RNN). The input of the RNN is a vector of the size of the state space, times the action space. The result of the last layer is passed through a softmax function to give the actual output of the network. This output is a vector of a probability distribution for all actions from which the control can select its next step by weighing the probabilities. Thus, once the control chooses to perform an action in a state, and the joint system either accepts or rejects this action, the RNN is updated to reflect the effects of taking that action. We use this latter solution as the underlying learner for the control agent, as it is the most unstructured solution, not modeling reward or goal estimations but only on the state and actions.
@@
As often the black box agent in the shared control system is actually a human or a cooperating agent [13, 14, 16, 21], we wish to leverage an additional communication channel, where the black box assumption is relaxed to allow the control to query the black box about the next best action it should take. Querying in such a manner to improve learning has been used in a variety of learning and inference problems, including active machine learning [4, 19, 24], diagnosis [10, 27], robotics [2, 30] and planning [5, 18, 26]. However, to the best of our knowledge, this work is the first to leverage active learning in shared control systems. We formalize this new communication channel by incorporating an oracle into the shared control system.
III The Query-based Shared Control Problem
@@
We start by formalizing the query-based shared control problem using a Multi-Agent Markov Decision Process (MA-MDP) that combines the state and action spaces of the black box agent and the control [3]. The key uniqueness of this model is in splitting the state space into two, according to the visibility of the agents’ state, from the learning agent’s perspective. This separation of the state space can also be viewed as an extension of Hidden-mode MDP into multiagent settings [8]. Each of the two spaces can capture an aggregated state of more than one agent in the environment and thus model systems with more than two agents. However, for brevity, in this paper we focus on the simplest case where the hidden part of the state space represents the state of a single black-box agent ($\mathbb{E}$), and the visible part represents the state space of the control ($\mathbb{C}$).
@@
  • $S^{\mathbb{C}}\times S^{\mathbb{E}}$ is a finite set of states such that $S^{\mathbb{C}}$ are the control’s states and $S^{\mathbb{E}}$ are the black box’s states.
  • $A^{\mathbb{C}}\times A^{\mathbb{E}}$ is a finite set of actions such that $A^{\mathbb{C}}$ are the control’s actions and $A^{\mathbb{E}}$ are the black box’s actions.
  • $T:S^{\mathbb{C}}\times S^{\mathbb{E}}\times A^{\mathbb{C}}\times A^{\mathbb{E}}\rightarrow S^{\mathbb{C}}\times S^{\mathbb{E}}$ is the transition of the shared system according to action taken by the control and the black box.
  • $R:S^{\mathbb{C}}\times S^{\mathbb{E}}\rightarrow\mathbb{R}$ is a function describing the reward of the agents for being at a specific joint state.
The objective of the shared control can be defined using $R$ to maximize different objectives, such as reaching some goal (as in shared driving, like example 1), maximizing the return (as in atari games, like the lunar lander from Section VI) or synchronizing the actions of the components (as in a dance performance, or in the automata from Section V-A1). In this paper, we keep the transition function deterministic, but it can be extended to stochastic transitions as well. The behavior of the transition function will be determined according to an Operation Protocol (OP), which defines the way that the actions of the agents will be combined.
@@
The proposed definition of a shared control system is very broad and can encompass a variety of shared control protocols. Different protocols can let one agent or both manipulate the joint operation or override an action. Alternatively, if the state and action spaces of the control and the black box are the same, a share control system can blend the chosen actions into some weighted sum.
Definition 2
An oracle is a function that gets a joint state and returns a set of actions $O:S^{\mathbb{C}}\times S^{\mathbb{E}}\rightarrow\mathcal{P}(A^{\mathbb{C}}\cup A^{\mathbb{E}})$.
We do not restrict the answer of the oracle to be an action that is executable by the control or the black box — the answer depends on the type of oracle used. E.g. a teacher oracle is familiar with the system as a whole and can provide the best action(s) the control should take to accomplish the shared objective over the course of an entire execution:
$O_{Teacher}(s^{\mathbb{C}},s^{\mathbb{E}})=\{\operatorname*{arg\,max}_{a\in A^{\mathbb{C}}}G(s^{\mathbb{C}},a)\}$
Where $G(s^{\mathbb{C}},a)$ can lead to the best potential return if the control takes action $a$ in $s^{\mathbb{C}}$. An expert oracle is one that is familiar with the black box’s behavior and can predict what is the next action that will maximize its reward.
$O_{Expert}(s^{\mathbb{C}},s^{\mathbb{E}})=\{\operatorname*{arg\,max}_{a\in A^{\mathbb{E}}}R(s^{\mathbb{E}},a)\}$
@@
IV Query Heuristics for a Shared Control
To answer the question of when to query the oracle, we propose three different heuristics, each inspired by different factors: The Entropy-based heuristic compares the Entropy from the learned policy’s suggestion and the expected Entropy from using the oracle; The Utility heuristic reasons about the probability of getting a successful interaction with a query, conditioned by our prior knowledge about successful transitions that state; The RL-based heuristic takes a Q-learning approach and chooses to query if the expected reward from querying at a state is higher than not querying.
IV-A Entropy Heuristic
The gist of this heuristic is to reason about the confidence of the learner about the information that was collected so far (or whether there’s a need to use an oracle). Information Gain (IG) is a measurement that can provide an answer to this question. It is a measure to estimate the confidence level of our learned policy as a predictor for the system’s next best action, using entropy estimation. Entropy is defined as the randomness or measuring the disorder of the information in our process [25]. It is also a machine learning metric that measures the unpredictability or impurity in the system. When entropy is fairly low, this heuristic provides a useful way to alert for the need for oracle queries. However, if the learned policy (e.g. using the RNN) has high confidence in a wrong action, the Entropy-based metric will only exacerbate this misconception, as seen in the empirical results in Section VI. Generally, given a discrete random variable $v$, which takes values in the alphabet $X$ and is distributed according to $p$:$X$→[0,1], entropy is defined as:
${\displaystyle\mathrm{H}(v):=-\sum_{x\in X}p(x)\log p(x)=\mathbf{E}[-\log p(X)]}$ (1)
The expected information gain from executing $q$ in state $S$ is the difference between the general entropy $\mathrm{H}(S)$ and the entropy given that $q$ was executed ($\mathrm{H}(S\mid q)$). Formally:
$\mathrm{IG}\left(\mathrm{S,q}\right)\mathrm{=H(S)-H(S\mid q)}$ (2)
When considering the IG with respect to an oracle, $S$ is the state of the joint system and $q$ is the query. Then, $H(S)$ is the entropy when acting in the environment without querying, and similarly $H(S\mid q)$ is the conditional entropy of $S$ given that we query the oracle. When the IG is bigger than a certain threshold, the system will use the assistance of an oracle to choose the next step.
Computing IG A challenging part of using the entropy heuristic to decide when to query is computing the IG. Given a shared control system $M$, $a^{\mathbb{C}}\in A^{\mathbb{C}}$ the action that was chosen by the control, $a^{\mathbb{E}}\in A^{\mathbb{E}}$ the action that was chosen by the black box, and $a^{o}\in A^{\mathbb{C}}\cup A^{\mathbb{E}}$ the action that was chosen by the oracle. The IG of the oracle for a state $s\in S^{\mathbb{C}}\times S^{\mathbb{E}}$ is the difference between querying or not:
${IG(s,querying)}=H(s)-H(s|querying){}$ (3)
The entropy H($s$) when not using the query is then given by:
${H(s)}=-\displaystyle\sum_{a^{\mathbb{C}}\in A^{\mathbb{C}}}p(s,a^{\mathbb{C}})log(p(s,a^{\mathbb{C}}))$ (4)
The value of $p(s,a^{\mathbb{C}})$ is given as the probability of $a^{\mathbb{C}}$ to be the action chosen by the RNN. Before the RNN is trained, these probabilities are initialized to be uniform. Going back to Equation 3, $H(s|querying)$ is harder to compute, as it is the entropy of the control when using the query at state $s$. To compute the overall entropy of the control for all possible oracle responses, we sum this value for all possible $a^{o}$, as given by equation 5:
@@ IV-A Entropy Heuristic
Intuitively, $H(s,a^{o})$ represents the entropy of the control’s choices when using the oracle, and receiving from it action $a^{o}$ in a state $s$. To compute it, we sum this value for each chosen action $a^{\mathbb{C}}$, given that the oracle’s choice was $a^{o}$, and multiply it by the probability of using the oracle:
$H($ $s,a^{o})=p(s,a^{o})H(s|a^{o})=$ (6)
$-p(s,a^{o})\sum_{a^{\mathbb{C}}\in A^{\mathbb{C}}}p(s,a^{\mathbb{C}}|a^{o})log(p(s,a^{\mathbb{C}}|a^{o}))$
As the probability of $a^{o}$ is not known in advance, it is estimated using the RNN as in Equation 4. However, since one cannot trust the oracle in states in which it is prone to mislead the control, it can bias the result of this computation to favor the wrong action due to false confidence. To avoid this bias, the next heuristic conditions the use of a query by the prior knowledge about the control’s past selections.
To conclude, given a state $s$ and a threshold $\beta_{ent}$, this heuristic chooses to query ($Oracle(s)$) or use its learned action ($Learned(s)$) according to the following condition:
@@
IV-B Utility Heuristic
This heuristic adds a preliminary step before choosing whether to query or not. When reaching a state $s$, it first checks the actions that are available for the control in $s$, and will choose to use the suggestion of the RNN if it gives the action a probability higher than some chosen threshold $\beta_{util}$.
@@ IV-B Utility Heuristic
IV-C Reinforcement Learning Heuristic
Reinforcement Learning (RL) is a type of machine learning in which an agent learns a policy by interacting with the environment and evaluating the outcomes of these interactions. Specifically, an RL problem is a tuple $\langle S,A,R\rangle$ where $S$ is a set of states, $A$ is a set of actions the agent can execute in the environment, and $R:S\times A\rightarrow\mathbb{R}$ is the reward for performing action $a$ at a state $s$ (we do not model the transition function as it is not known to the control). In the shared control learning context, the state space used in this heuristic is the states of the visible control $S^{\mathbb{C}}$, the action space consists of two actions: to query and not to query, and the reward is defined to some positive reward if after a choice whether to query or not to query. Notice that this MDP is different from the MA-MDP used to model the shared control system as a whole, where the key difference is that here the decision the agents needs to make is only about querying or not. The RL heuristic is based on the Q-learning algorithm [28], where the agent keeps an estimation of the value of executing an action in a state using a value function. This function is called a Q-function and it is defined as $Q:S\rightarrow A$. The heuristic chooses to query if the value of querying is higher than not, according to the Q-function.
@@ IV-C Reinforcement Learning Heuristic
V Empirical Evaluation
The underlying learning agent used in this work is the one presented in Iosti et al. [12]. However, as mentioned earlier, it can be replaced by any other learning agent that can process the states and actions of the system and output a recommendation for the next action of the control. The empirical evaluation in this paper consists of two domains: an automata-based system and an Atari-based lunar lander simulation [7].
V-A Domains
V-A1 Automata-based Shared Control
We follow the definition of an automaton by Iosti et al. [12]. These automata are carefully chosen to challenge the learning of the control policy, due to sparse or delayed rewards.
An automaton A is a tuple $\Omega=\langle S,\iota,A,\delta\rangle$, where $S$ is a finite set of states, $\iota\in S$ it the initial state of the automaton, $A$ is a finite set of actions, and $\delta:(S\times A)\rightarrow S\cup\{\perp\}$ is a partial transition function, where $\perp$ stands for an undefined transition. For brevity, we use $a(s)$ to define the transition $\delta(a,s)$. Denoting $en(s)=\{a\mid a\in A\wedge\delta(s,a)\neq\perp\}$, i.e., $en(s)$ is the set of actions enabled at the state $s$. We assume that for each $s\in S$, the set of possible actions is not empty, $en(s)\neq\emptyset$. A shared control system consists of two automata, a black-box ($\Omega^{\mathbb{E}}$) and the control ($\Omega^{\mathbb{C}}$).
As this is a black-box environment, the control automaton is visible to the learner, but the black box automaton is not. The operation protocol for the two automata is an agreement operator, which is defined as: